OpenStudy (anonymous):

A baker made 20 pies. A Boy Scout troop buys one–fourth of his pies, a preschool teacher buys one–third of his pies, and a caterer buys one–sixth of his pies. How many pies does the baker have left?

OpenStudy (anonymous):

Okay so \[20\times \frac{ 1 }{ 4 }=5\]So now the baker has 15 pies\[15\times \frac{ 1 }{ 3}=5\]Now the baker has 10 pies \[10\times \frac{ 1 }{ 6 }=\frac{ 5 }{ 3}\]Also known as 1 and 2/3. The baker should have 8 and 1/3 pies left

OpenStudy (anonymous):

Thanks